Skip to main content

Note 1 to 6

Combination and Permutation​

P(n, r) = n! / (n - r)!
C(n, r) = P(n, r) / r! = n! / (r!(n - r)!)

Star and bar 隔板法​

https://zh.wikipedia.org/wiki/%E9%9A%94%E6%9D%BF%E6%B3%95

n object into k box
x1 + x2 + ... xk = n

Not allow empty box (at least 1)
C(n - 1, k - 1)

Allow empty box (at least 0)
C(n + k - 1, k - 1)

At least 2
C(n - k - 1, k - 1)

At least 3
C(n - 2 * k - 1, k - 1)

At least m
C(n + (1 - m) * k - 1, k - 1)

x1 + x2 + ... xk = n
where xi >= mi (at least mi)
C(n + sum of (1 - mi) - 1, k - 1)

Example:
x1 + x2 + x3 + x4 = 20
where x1 >= 0, x2 >= 1, x3 >= 2, x4 >= 3
How many integer solution?

C(20 + ((1 - 0) + (1 - 1) + (1 - 2) + (1 - 3)) - 1, 4 - 1)
= C(20 + (1 + 0 - 1 - 2) - 1, 4 - 1)
= 680

插空法​

https://zh.wikipedia.org/wiki/%E6%8F%92%E7%A9%BA%E6%B3%95

Problem of points / Unfinished Game / 賭金分配問題 The Problem of Division of the Stakes /​

money * win round of someone / all possible outcomes of remaining rounds
list of all the outcome and count

A remain x round to win
B remain y round to win
Max total remaining rounds = (x - 1) + (y - 1) + 1 = x + y - 1

Xi remain wi round to win, where i from 1 to n
Max total remaining rounds = sum of (wi - 1) + 1 = sum of wi - n + 1

(Out of Syllabus)
// https://en.wikipedia.org/wiki/Problem_of_points
ratio of A : ratio of B
ratio of A = Sum of k = 0 to s - 1 ( C(r + s - 1, k) )
ratio of B = Sum of k = s to r + s - 1 ( C(r + s - 1, k) )

Two Children Problem / Boy or Girl paradox​

A Family with two children. One of the children is a boy, what is the probability that both children are boys?

BB BG GB
1 / 3


A Family with two children. One of the children is a girl both in pm, what is the probability that other child is boy both in am?

BB BG GB GG
Ba Ba Ba Ga Ga Ba Ga Ga
Ba Bp Ba Gp Ga Bp Ga Gp
Bp Ba Bp Ga Gp Ba Gp Ga
Bp Bp Bp Gp Gp Bp Gp Gp

(Filter by one of children is girl both in pm)

BB BG GB GG

Ba Gp Ga Gp
Gp Ba Gp Ga
Bp Gp Gp Bp Gp Gp

(Ba Gp, Gp Ba match)
2 / 7

// https://en.wikipedia.org/wiki/Boy_or_Girl_paradox

Birthday Problem​

n people in the same room, what is the probability of two people born on same day

Probabilty of even one not same birthday:
1 * (1 - 1 / 365) * (1 - 2 / 365) * (1 - (n - 1) / 365)
= 365 / 365 * 364 / 365 * 363 / 365 ... (365 - n + 1) / 365
= 365! / (365^n (365 - n)!)

p(n) = 1 - 365! / (365^n (365 - n)!)

False Positive​

Test Always positive if have disease
Probability of false positive = fp
Probability of have disease = d

Positive Negative
Disease d * 1 = d 0
No disease d' * fp Not given

P(disease | positive) = d / (d + d' * fp)

Three Axioms of Probability​

1. P(A) >= 0 for any event A
2. P(S) = 1, where S is sample space
3. P(A1 U A2 U A3 ...) = P(A1) + P(A2) + P(A3) + ...

Notation​

P(A U B) = P(A or B)
P(A ∩ B) = P(AB) = P(A and B)
Usually P(A ∩ B) will write as P(AB) in following

Basic​

P(A) <= 1
P(A') = 1 - P(A)

P(A - B) = P(AB')
P(A − B) = P(A) − P(AB)

Union​

P(A U B) = P(A) + P(B) - P(AB)

If A, B is disjoint
P(A U B) = P(A) + P(B)

Intercept​

P(AB) = P(A) P(B)

If A, B is disjoint
P(AB) = 0

If A, B is dependent
P(AB) = P(A) P(B | A) = P(B) P(A | B)

If A, B is independent
P(AB) = P(A) P(B)
P(A1 A2 A3 ...) = P(A1) P(A2) P(A3) ...

Disjoint event​

Disjoint mean A and B cannot occur at the same time
Disjoint != Independent

AB = ∅
P(AB) = P(∅) = 0
P(A U B) = P(A) + P(B)
P(A1 U A2 U A3 ...) = P(A1) + P(A2) + P(A3) + ...

Dependent event​

Dependent mean happen A or B will affect probability of other

P(AB) = P(A) P(B | A) = P(B) P(A | B)

Independent event​

Independent mean happen A or B not affect probability of other
Independent != Disjoint

P(AB) = P(A) P(B)
P(A | B) = P(A)

A and B' are also independent
A' and B are also independent
A' and B' are also independent

Use P(AB) != P(A) P(B) to prove A, B not Independent event

E(XY) = E(X) E(Y)
Var(X + Y) = Var(X) + Var(Y)

Cumulative distribution function (CDF)​

F(X) = P(X <= x)

Binomial Distribution​

X ~ B(n, p)
P(X = k) = C(n, k) p^k (1 - p)^(n - k)

E(X) = np
Var(X) = np(1 - p)

Expectation Mean​

E(X) = sum of  x P(X = x)
E(X) = sum of (1 - P(X <= x)) = sum of P(X > x)
E(f(x)) = sum of f(x) P(X = x)

E(X + Y) = E(X) + E(Y)
E(aX) = aE(x)

If Independent
E(XY) = E(X)E(Y)

If X ~ B(n, p)
E(X) = np

Variance​

Var(X) = E([X - E(X)]^2)
Var(X) = E(X^2) - E(X)^2

Var(X + a) = Var(X)
Var(aX) = a^2 Var(X)
Var(-X) = Var(X)

If Independent
Var(X + Y) = Var(X) + Var(Y)

If X ~ B(n, p)
Var(X) = np(1 - p)

Standard Deviation​

SD(X) = sqrt(Var(X))

SD(aX + b) = |a|SD(X)

Markov’s Inequality​

P(X >= a) <= E(X) / a  if X >= 0, a > 0

Chebyshev’s Inequality​

P(|X - E(X)| >= k * SD(x)) <= 1 / k^2
P(|X - E(X)| < k * SD(x)) >= 1 - 1 / k^2

Bernoulli’s Utility​

utility = log( wealth )
utility change of gain = log( wealth + gain ) – log( wealth )